ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 16N80 Zand, H, Open University, England, and Edmund Puczylowski, University of Warsaw, Poland Duality in radicals of associative rings Radical theory, initiated by Kurosh, has played an important role in the theory of rings. In this talk a duality between certain radicals will be introduced and new results will be presented.